kayl612:

The equation of a circle is x2 + y2 + Cx + Dy = E. If the radius of the circle is decreased without changing the coordinates of the center point, how are the coefficients C, D, and E affected?

surjithayer:

@surjithayer wrote:
\[x^2+y^2+Cx+Dy=E\] \[x^2+Cx+(\frac{ C }{ 2 })^2+y^2+Dy+(\frac{ D }{ 2 })^2=E+(\frac{ C }{ 2 })^2+(\frac{ D }{ 2 })^2\] \[(x+\frac{ C }{ 2 })^2+(y+\frac{ D }{ 2 })^2=(\sqrt{E+\frac{ C^2 }{ 4 }+\frac{ D^2 }{ 4 }})^2\] center\[=(-\frac{C}{2},-\frac{D}{2})\] \[radius=\sqrt{E+\frac{ C^2 }{ 4 }+\frac{ D }{ 4 }}\] as C and D are fixed. so for radius to change only E changes.